OpenStudy (anonymous):

please help meee solve for x: (x-4)/x+2y=W/3 What I did was cross multiply and got 3(x-4)=W(x+2y).

OpenStudy (anonymous):

after distributing I got 3x-12=wx+2wy. Is that correct?

OpenStudy (anonymous):

3x-12=wx+2yw 3x-wx=12+2yw factor out the x on the left x(3-w)=12+2yw divide both sides by 3-w x=(12+2yw)/(3-w)

OpenStudy (anonymous):

yay thanks! I would like to be your fan :)